1. A researcher reports an outcome of one-factor ANOVA for independent samples as “the results showed a significant difference among treatments, F(3, 64) = 5.42, p < .01.” How many treatment groups were compared in this analysis?

  • A. 4

  • B. 5

  • C. 3

  • D. 1

  • E. 2

2. In one-factor AOVA, which of the following is expected If there is no significant effect of experimental treatments?

  • A. MSbetween should be greater than MSwithin

  • B. MSbetween should be about the same size as MSwithin

  • C. MSwithin should be greater than MSbetween

  • D. SSbetween should be about the same size as SStotal

3. A researcher uses a one-factor ANOVA to examine differences among five treatment means with a separate sample of n = 10 participants in each treatment. What are the df values for the F-ratio in this analysis?

  • A. 4, 50

  • B. 4, 45

  • C. 5, 50

  • D. 5, 40

  • E. 5, 10
Answer #1

1. Number of treatment groups = Treatment df + 1 = 3 + 1 = 4

Option A is correct.

2. MSbetween should be about the same size as MSwithin

Option B is correct.

3. Treatment df = 5 - 1 = 4

Error df = 5*10 - 5 = 45

Option B is correct.